# Expressive Power of Graph Neural Networks for (Mixed-Integer) Quadratic Programs

### v1

@@ 2. Preliminaries
We define three mappings for LCQPs.
  • Feasibility mapping: $\Phi_{\textup{feas}}(G_{\textup{LCQP}})=1$ if the LCQP problem associated to $G_{\textup{LCQP}}$ is feasible and $\Phi_{\textup{feas}}(G_{\textup{LCQP}})=0$ if it is infeasible.
  • Optimal objective value mapping: $\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}\cup\{\pm\infty\}$ computes the optimal objective value of the LCQP problem associated to $G_{\textup{LCQP}}$. $\Phi_{\textup{obj}}(G_{\textup{LCQP}})=+\infty$ means the problem is infeasible and $\Phi_{\textup{obj}}(G_{\textup{LCQP}})=-\infty$ means the problem is unbounded.
  • Optimal solution mapping: For an feasible and bounded LCQP problem $G_{\textup{LCQP}}$, i.e., $\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}$, there must exist an optimal solution [eaves1971quadratic] though it might not be unique. However, the optimal solution with the smallest $\ell_{2}$-norm must be unique if $Q\succeq 0$ by similar argument as in [chen2022representing-lp]*Remark 2.2 and we define it as $\Phi_{\textup{sol}}(G_{\textup{LCQP}})$.
@@
3. Unversal approximation of GNNs for LCQPs
This section presents our main theoretical results for the expressive power of GNNs for representing properties of linearly constrained quadratic programs (LCQPs). In particular, we show that for any LCQP data distribution, there always exists a GNN that can predict LCQP properties, in the sense of universally approximating target mappings in Definition 2.3, within some given error tolerance. Such results answer the question (1.1) positively. We state the assumption required for our main theorems.
@@ 3. Unversal approximation of GNNs for LCQPs
[…]
The assumption of Borel regularity is generally satisfied for most data distributions in practice, including discrete distributions, gaussian distributions, etc.
We first state the universal approximation result of MP-GNNs for representing the feasibility of LCQP directly implied by the previous work [chen2022representing-lp] on linear programs, as feasibility is solely determined by the constraints, independent of the objective function, and all LCQP constraints are linear.
Theorem 3.2 ([chen2022representing-lp]*Theorem 3.2).
For any probability measure $\mathbb{P}$ on $\mathcal{G}^{m,n}_{\textup{LCQP}}$ satisfying Assumption 3.1 and any $\epsilon>0$, there exists $F\in\mathcal{F}_{\textup{LCQP}}$ such that $\mathbb{I}_{F(G_{\textup{LCQP}})>\frac{1}{2}}$ can act as a classifier for LCQP-feasibility, with an error of up to $\epsilon$:
$\mathbb{P}\big{[}\mathbb{I}_{F(G_{\textup{LCQP}})>\frac{1}{2}}\neq\Phi_{\textup{feas}}(G_{\textup{LCQP}})\big{]}<\epsilon,$
@@
[…]
@@
Let $\mathbb{P}$ be a probability measure on $\mathcal{G}^{m,n}_{\textup{LCQP}}$ satisfying Assumption 3.1.
  1. (1)
    For any $\epsilon>0$, there exists $F_{1}\in\mathcal{F}_{\textup{LCQP}}$ such that
    $\mathbb{P}\big{[}\mathbb{I}_{F_{1}(G_{\textup{LCQP}})>\frac{1}{2}}\neq\mathbb{I}_{\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}}\big{]}<\epsilon.$
  2. (2)
    If $\mathbb{P}[\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}]=1$, then for any $\epsilon,\delta>0$, there exists $F_{2}\in\mathcal{F}_{\textup{LCQP}}$ such that
    $\mathbb{P}\left[|F_{2}(G_{\textup{LCQP}})-\Phi_{\textup{obj}}(G_{\textup{LCQP}})|>\delta\right]<\epsilon.$
Our last theorem for LCQP is that GNN can approximate the optimal solution map $\Phi_{\textup{sol}}$ that returns the optimal solution with the smallest $\ell_{2}$-norm of feasible and bounded LCQP problems.
@@
Let $\mathbb{P}$ be a probability measure on $\mathcal{G}^{m,n}_{\textup{LCQP}}$ satisfying Assumption 3.1 and assume that $\mathbb{P}[\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}]=1$. For any $\epsilon,\delta>0$, there exists $F_{W}\in\mathcal{F}_{\textup{LCQP}}^{W}$ such that
@@
The proofs of Theorem 3.3 and Theorem 3.4 will be presented in Appendix A. We briefly describe the main idea here. The Stone-Weierstrass theorem and its variants are a powerful tool for proving universal-approximation-type results. Recall that the classic version of the Stone-Weierstrass theorem states that under some assumptions, a function class $\mathcal{F}$ can uniformly approximate every continuous function if and only if it separates points, i.e., for any $x\neq x^{\prime}$, one has $F(x)\neq F(x^{\prime})$ for some $F\in\mathcal{F}$. Otherwise, we say $x$ and $x^{\prime}$ are indistinguishable by any $F\in\mathcal{F}$. Therefore, the key component in the proof is to establish some separation results in the sense that two LCQP-graphs with different optimal objective values (or different optimal solutions with the smallest $\ell_{2}$-norm) must be distinguished by some GNN in the class $\mathcal{F}_{\textup{LCQP}}$ (or $\mathcal{F}_{\textup{LCQP}}^{W}$). It is first established in [xu2019powerful] that the separation power [^2] Given two sets of functions, $\mathcal{F}$ and $\mathcal{F}^{\prime}$, both defined over the same domain $X$, if $\mathcal{F}$ separating points $x$ and $x^{\prime}$ implies that $\mathcal{F}^{\prime}$ also separates $x$ and $x^{\prime}$ for any $x,x^{\prime}\in X$, then the separation power of $\mathcal{F}^{\prime}$ is considered to be stronger than or at least equal to that of $\mathcal{F}$. of message-passing GNNs is equivalent to the Weisfeiler-Lehman (WL) test [weisfeiler1968reduction], a classical algorithm for the graph isomorphism problem, which is further developed in many recently works, see e.g. [azizian2020expressive, geerts2022expressiveness, chen2022representing-lp]. We show that, any two LCQP-graphs that are indistinguishable by the WL test, or equivalently by all message-passing GNNs, even if they are not isomorphic, some of their structures must be identical, which guarantees that they must have identical optimal objective value and identical optimal solution with the smallest $\ell_{2}$-norm.
@@ 4. The capacity of GNNs for MI-LCQPs
In this section, we discuss the expressive power of GNNs for mixed-integer linearly constrained quadratic programs (MI-LCQPs), for which the general form is almost the same as (2.1) except that some entries of $x$ are constrained to be integers: $x_{j}\in\mathbb{Z},\ \forall~{}j\in I$, where $I\subset\{1,2,\dots,n\}$ collects the indices of all integer variables. Before proceeding, we extend LCQP-graphs and the corresponding GNNs and target mappings to the MI-LCQP setting.
MI-LCQP-graph is modified from the LCQP-graph (Section 2 and Figure 1) by adding a new entry to the feature of each variable node $j\in W$. The new feature is $w_{j}=(c_{j},l_{j},u_{j},\delta_{I}(j))$ where $\delta_{I}(j)=1$ if $j\in I$ and $\delta_{I}(j)=0$ otherwise. We use $\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ to denote the collection of all MI-LCQP-graphs with $m$ constraints, $n$ variables, and symmetric and positive semi-definite $Q$.
GNNs for MI-LCQP-graphs are constructed following the same mechanism as for LCQP-graphs, with the difference that the message-passing layer is modified as
  • $s_{i}^{l}=f_{l}^{V}\big{(}s_{i}^{l-1},\sum_{j\in\mathcal{N}_{i}^{W}}g_{l}^{W}(t_{j}^{l-1},A_{ij})\big{)}$ for $i\in V$, and
  • $t_{j}^{l}=f_{l}^{W}\big{(}t_{j}^{l-1},\sum_{i\in\mathcal{N}_{j}^{V}}g_{l}^{V}(s_{i}^{l-1},A_{ij}),\sum_{j^{\prime}\in\mathcal{N}_{j}^{W}}g_{l}^{Q}(t_{j^{\prime}}^{l-1},Q_{jj^{\prime}})\big{)}$ for $j\in W$,
where $\mathcal{N}_{i}^{W}=\{j\in W:A_{ij}\neq 0\}$, $\mathcal{N}_{j}^{V}=\{j\in V:A_{ij}\neq 0\}$, and $\mathcal{N}_{j}^{W}=\{j^{\prime}\in W:Q_{jj^{\prime}}\neq 0\}$ are the sets of neighbors. We use $\mathcal{F}_{\textup{MI-LCQP}}$ and $\mathcal{F}_{\textup{MI-LCQP}}^{W}$ to denote the GNN classes for MI-LCQP-graphs with graph-level and node-level output, respectively.
Target mappings for MI-LCQPs considered in this section are also similar to those in Definition 2.3. In particular, the feasibility mapping $\Phi_{\textup{feas}}$ and the optimal objective value mapping $\Phi_{\textup{obj}}$ are defined in the same way as in Definition 2.3, while the optimal solution mapping $\Phi_{\textup{sol}}$ can only be defined on a subset of the class of feasible and bounded MI-LCQPs, which will be discussed in Section 4.2.

### v2

@@ 2. Preliminaries
We define:
  • Feasibility mapping: $\Phi_{\textup{feas}}(G_{\textup{LCQP}})=1$ if $G_{\textup{LCQP}}$ is feasible and $\Phi_{\textup{feas}}(G_{\textup{LCQP}})=0$ if it is infeasible.
  • Optimal objective mapping: $\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}\cup\{\pm\infty\}$ computes the optimal objective value of $G_{\textup{LCQP}}$. $\Phi_{\textup{obj}}(G_{\textup{LCQP}})=+\infty$ means the problem is infeasible and $\Phi_{\textup{obj}}(G_{\textup{LCQP}})=$-$\infty$ means unboundedness.
  • Optimal solution mapping: For a feasible and bounded LCQP problem (i.e., $\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}$), an optimal solution exists though it might not be unique. However, the optimal solution with the smallest $\ell_{2}$-norm must be unique if $Q\succeq 0$, i.e., $Q$ is positive semi-definite, and we define it as $\Phi_{\textup{sol}}(G_{\textup{LCQP}})$.
@@
3. Universal approximation for convex LCQPs
This section presents our main theoretical results for the expressive power of GNNs for representing properties of LCQPs. In particular, we show that for any convex LCQP data distribution, there always is a GNN that can predict LCQP properties, in the sense of universally approximating target mappings in Definition 2.3, within a given error tolerance. Although it is known in the previous literature that there exists some continuous function that cannot be approximated by GNNs with arbitrarily small error, see e.g., [xu2019powerful, azizian2020expressive, geerts2022expressiveness], our results in this section indicate that approximating the target mappings of LCQPs do not suffer from this limitation.
@@ 3. Universal approximation for convex LCQPs
$\mathbb{P}$ is a Borel regular probability measure defined on the space of LCQP-graphs $\mathcal{G}^{m,n}_{\textup{LCQP}}$.
The assumption of Borel regularity is generally satisfied for most data distributions in practice.
Theorem 3.2.
For any $\mathbb{P}$ satisfying Assumption 3.1 and any $\epsilon>0$, there exists $F\in\mathcal{F}_{\textup{LCQP}}$ such that $\mathbb{I}_{F(G_{\textup{LCQP}})>\frac{1}{2}}$ acts as a classifier for LCQP-feasibility, with an error of up to $\epsilon$:
$\mathbb{P}\left[\mathbb{I}_{F(G_{\textup{LCQP}})>\frac{1}{2}}\neq\Phi_{\textup{feas}}(G_{\textup{LCQP}})\right]<\epsilon,$
@@
This result suggests that a GNN is a universal classifier for LCQP feasibility: for any data distribution of LCQPs satisfying Assumption 3.1, there exists a GNN that can classify LCQP feasibility with arbitrarily high accuracy. This is a natural extension of the feasibility classification for linear programs [chen2022representing-lp], as feasibility is solely determined by the constraints, independent of the objective function, and all LCQP constraints are linear.
However, using GNNs to predict the optimal objective value or an optimal solution is highly non-trivial due to the nonlinear term $x^{\top}Qx$. Fortunately, when restricting LCQPs to convex cases, GNNs can universally represent the optimal objective value and an optimal solution for these LCQPs.
@@
For any $\mathbb{P}$ satisfying Assumption 3.1 with $\mathbb{P}[Q\succeq 0]=1$ ($Q$ is positive semidefinite almost surely), and for any $\epsilon>0$, there exists $F_{1}\in\mathcal{F}_{\textup{LCQP}}$ such that
(3.1) $\mathbb{P}\left[\mathbb{I}_{F_{1}(G_{\textup{LCQP}})>\frac{1}{2}}\neq\mathbb{I}_{\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}}\right]<\epsilon.$
Addtitionally, if $\mathbb{P}[\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}]=1$, then for any $\epsilon,\delta>0$, there exists $F_{2}\in\mathcal{F}_{\textup{LCQP}}$ such that
(3.2) $\mathbb{P}\left[|F_{2}(G_{\textup{LCQP}})-\Phi_{\textup{obj}}(G_{\textup{LCQP}})|>\delta\right]<\epsilon.$
This theorem indicates that GNNs can approximate the optimal objective value mapping $\Phi_{\textup{obj}}$ very well in two senses: (I) GNN can predict whether the optimal objective value is a real number or $\pm\infty$, i.e., whether the LCQP problem is feasible and bounded or not. (II) For a data distribution over feasible and bounded LCQP problems, GNN can approximate the optimal objective. Finally, we prove that GNN can approximate the optimal solution map $\Phi_{\textup{sol}}$.
@@
For any $\mathbb{P}$ satisfying Assumption 3.1 with $\mathbb{P}[Q\succeq 0]=\mathbb{P}[\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}]=1$, and for any $\epsilon,\delta>0$, there exists $F_{W}\in\mathcal{F}_{\textup{LCQP}}^{W}$ such that
@@
The detailed proofs of Theorems 3.3 and 3.4 will be presented in Appendix A. We briefly describe the main idea here. The Stone-Weierstrass theorem and its variants are a powerful tool for proving universal-approximation-type results. Recall that the classic version of the Stone-Weierstrass theorem states that under some assumptions, a function class $\mathcal{F}$ can uniformly approximate every continuous function if and only if it separates points, i.e., for any $x\neq x^{\prime}$, one has $F(x)\neq F(x^{\prime})$ for some $F\in\mathcal{F}$. Otherwise, we say $x$ and $x^{\prime}$ are indistinguishable by any $F\in\mathcal{F}$. Therefore, the key component in the proof is to establish some separation results in the sense that two LCQP-graphs with different optimal objective values (or different optimal solutions with the smallest $\ell_{2}$-norm) must be distinguished by some GNN in the class $\mathcal{F}_{\textup{LCQP}}$ (or $\mathcal{F}_{\textup{LCQP}}^{W}$). It is shown in [xu2019powerful, azizian2020expressive, geerts2022expressiveness] that the separation power [^2] Given two sets of functions, $\mathcal{F}$ and $\mathcal{F}^{\prime}$, both defined over domain $X$, if $\mathcal{F}$ separating points $x$ and $x^{\prime}$ implies that $\mathcal{F}^{\prime}$ also separates $x$ and $x^{\prime}$ for any $x,x^{\prime}\in X$, then the separation power of $\mathcal{F}^{\prime}$ is considered to be stronger than or equal to that of $\mathcal{F}$. of GNNs is equivalent to the Weisfeiler-Lehman (WL) test [weisfeiler1968reduction], a classical algorithm for the graph isomorphism problem. We show that, any two LCQP-graphs that are indistinguishable by the WL test, or equivalently by all GNNs, even if they are not isomorphic, they must have identical optimal objective value and identical optimal solution with the smallest $\ell_{2}$-norm.
An illustrative example
We use the two LCQPs, (3.3) and (3.4), each with 7 variables and 8 constraints, to illustrate our findings. All variable nodes share the attribute $w_{j}=(1,0,3)$ for $1\leq j\leq 7$, which represents an objective coefficient of $c_{j}=1$, lower bound $l_{j}=0$, upper bound $u_{j}=3$. We refer to these nodes as “red nodes.” The first seven constraint nodes $v_{i}$ (for $1\leq i\leq 7$) are assigned the same attribute, $v_{i}=(0,=)$, which we label as “blue nodes”. The eighth constraint node $v_{8}$ is unique, with the attribute $v_{8}=(6,=)$, and is called the “brown node.” Any red node is connected to a blue node with weight $A_{ij}=1$ (solid lines), another blue node with weight $A_{ij}=-1$ (dashed lines), the brown node with weight $A_{ij}=1$ (green lines), and all seven red nodes with $Q_{jj^{\prime}}=1$ (brown curves).
(3.3) $\begin{split}\min_{x\in\mathbb{R}^{7}}&~~\frac{1}{2}x^{\top}\mathbf{1}\mathbf{1}^{\top}x+\mathbf{1}^{\top}x,\\ \text{s.t.}&~~x_{1}-x_{2}=0,\ x_{2}-x_{1}=0,\\ &~~x_{3}-x_{4}=0,\ x_{4}-x_{5}=0,\\ &~~x_{5}-x_{6}=0,\ x_{6}-x_{7}=0,\ x_{7}-x_{3}=0,\\ &~~x_{1}+x_{2}+x_{3}+x_{4}+x_{5}+x_{6}+x_{7}=6\\ &~~0\leq x_{j}\leq 3,\ x_{j}\in\mathbb{Z},\ \forall~j\in\{1,2,\dots,7\}.\end{split}$
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$
(3.4) $\begin{split}\min_{x\in\mathbb{R}^{7}}&~~\frac{1}{2}x^{\top}\mathbf{1}\mathbf{1}^{\top}x+\mathbf{1}^{\top}x,\\ \text{s.t.}&~~x_{1}-x_{2}=0,\ x_{2}-x_{3}=0,\ x_{3}-x_{1}=0,\\ &~~x_{4}-x_{5}=0,\ x_{5}-x_{6}=0,\\ &~~x_{6}-x_{7}=0,\ x_{7}-x_{4}=0,\\ &~~x_{1}+x_{2}+x_{3}+x_{4}+x_{5}+x_{6}+x_{7}=6\\ &~~0\leq x_{j}\leq 3,\ x_{j}\in\mathbb{Z},\ \forall~j\in\{1,2,\dots,7\}.\end{split}$
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$
Figure 2. An illustrative example for Theorems 3.3 and 3.4
(Observation I). The two LCQPs are not equivalent, and their graph representations are not isomorphic. Both LCQPs are feasible and bounded, with identical optimal objective values, as $\mathbf{1}^{\top}x=6$ results in $\frac{1}{2}x^{\top}\mathbf{1}\mathbf{1}^{\top}x+\mathbf{1}^{\top}x=24$. However, they are not equivalent because their solution sets differ. Specifically, (3,3,0,0,0,0,0){(3,3,0,0,0,0,0)} is a solution to (3.3) but not to (3.4), while (2,2,2,0,0,0,0){(2,2,2,0,0,0,0)} is a solution to (3.4) but not to (3.3). Furthermore, their graph representations differ: in the first graph, the 7 variables form two groups—one with 2 nodes and the other with 5—where nodes in each group are connected cyclically. In the second graph, the variables form two groups, but one with 3 nodes and the other with 4.
(Observation II). The two graphs cannot be distinguished by any GNNs, even after multiple rounds of message passing. Initially, both graphs are indistinguishable as they share identical node attributes: seven red nodes, seven blue nodes, and one brown node. During message passing, each red node receives identical updates due to symmetric connections: one blue node with weight $A_{ij}=1$, another with $A_{ij}=-1$, the brown node with $A_{ij}=1$, and all seven red nodes with $Q_{jj^{\prime}}=1$. As a result, the red node’s attribute is updated as (an informal but illustrative equation):
$t_{j}^{l}=f_{l}^{w}\Big(\textup{red node},g_{l}^{V}(\textup{blue node},1)+g_{l}^{V}(\textup{blue node},-1)$
$+g_{l}^{V}(\textup{brown node},1),~7\cdot g_{l}^{Q}(\textup{red node},1)\Big).$
After the update, all red nodes $t_{j}^{l}(1\leq j\leq 7)$ in both graphs retain identical attributes and are still indistinguishable. The same applies to the blue and brown nodes. Therefore, regardless of how many message-passing rounds occur, both graphs will still have seven red nodes, seven blue nodes, and one brown node. This holds for any parameterized mappings used in GNNs ($f_{l}^{V}$, $f_{l}^{W}$, $g_{l}^{V}$, $g_{l}^{W}$, and $g_{l}^{Q}$), meaning no GNN can differentiate between the two instances.
(Observation III). They share the common optimal solution with the smallest $\ell_{2}$-norm: $(\frac{6}{7},\frac{6}{7},\frac{6}{7},\frac{6}{7},\frac{6}{7},\frac{6}{7},\frac{6}{7})$, which follows from $\mathbf{1}^{\top}x=6$ and the Cauchy-Schwarz inequality.
Note that these observations are not mere coincidences; we have established this conclusion in general for LCQPs (see Definition A.1, Theorem A.2, and Theorem A.3).
Additionally, similar results can be extended to QCQPs under certain assumptions, though it is beyond the main focus of this paper. The details are deferred to Appendix E.
A nonconvex counterexample
At the end of this section, we comment that Theorems 3.3 and 3.4 do not hold if the convexity assumption ($\mathbb{P}[Q\succeq 0]=1$) is removed. In particular, consider a convex LCQP
$\min~~\frac{1}{2}\begin{pmatrix}x_{1}&x_{2}\end{pmatrix}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix},\quad\text{s.t.}~-1\leq x_{1},x_{2}\leq 1,$
and a nonconvex LCQP
$\min~~\frac{1}{2}\begin{pmatrix}x_{1}&x_{2}\end{pmatrix}\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix},\quad\text{s.t.}~-1\leq x_{1},x_{2}\leq 1.$
Following the aforementioned argument, one can verify that these two LCQP instances are indistinguishable by any GNNs. On the other hand, the first convex LCQP has an optimal objective being 0 and the optimal solution being $(0,0)$, while the second nonconvex LCQP has an optimal objective being 1-1 and optimal solutions being (1,1)(1,-1) and (1,1)(-1,1). This indicates that the conclusions in Theorems 3.3 and 3.4 do not hold true for the data distribution that is supported on aforementioned instances with equal probability, since a GNN making accurate prediction on one instance must fail on the other.
@@ 4. The capacity of GNNs for MI-LCQPs
In this section, we turn to mixed-integer linearly constrained quadratic program (MI-LCQP), which is almost the same as LCQP (2.1) except that some entries of $x$ are constrained to be integers: $x_{j}\in\mathbb{Z},\ \forall~j\in I$, where $I\subset\{1,2,\dots,n\}$ collects the indices of all integer variables.
MI-LCQP-graph is modified from the LCQP-graph by adding a new entry to the feature of each variable node $j\in W$. The new feature is $w_{j}=(c_{j},l_{j},u_{j},\delta_{I}(j))$ where $\delta_{I}(j)=1$ if $j\in I$ and $\delta_{I}(j)=0$ otherwise. We use $\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ to denote the collection of all MI-LCQP-graphs.
GNNs for MI-LCQP-graphs are constructed using the same framework as those for LCQP-graphs, differing only in the input feature $w_{j}$ as defined above. We use $\mathcal{F}_{\textup{MI-LCQP}}$ and $\mathcal{F}_{\textup{MI-LCQP}}^{W}$ to denote the GNN classes for MI-LCQP-graphs with graph-level and node-level output, respectively.
Target mappings for MI-LCQPs also similar to those in Definition 2.3. In particular, $\Phi_{\textup{feas}}$ and $\Phi_{\textup{obj}}$ are defined as in Definition 2.3, while the optimal solution mapping $\Phi_{\textup{sol}}$ can only be defined on a subset of the class of feasible and bounded MI-LCQPs, as discussed in Appendix C.